OpenStudy (anonymous):

Find the angle between the given vectors to the nearest tenth of a degree. u = <-5, -4>, v = <-4, -3>

jimthompson5910 (jim_thompson5910):

here's a page with the formula and an example http://www.vitutor.com/geometry/vec/angle_vectors.html

OpenStudy (anonymous):

Use \[ \cos \theta = \frac{\mathbf{u}\cdot \mathbf v}{\|\mathbf u\|\|\mathbf v\|} \]

OpenStudy (anonymous):

@jim_thomson5190 im lost I don't see how they got the answer

jimthompson5910 (jim_thompson5910):

where are you stuck

OpenStudy (anonymous):

I work the problem out on the ex and I dont see how they got 45

jimthompson5910 (jim_thompson5910):

do you see how they got cos(alpha) = sqrt(2)/2 ??

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

well you want the value of alpha, not cos(alpha)

jimthompson5910 (jim_thompson5910):

so you must apply the arccosine to both sides

jimthompson5910 (jim_thompson5910):

that's how they got 45 degrees arccos(sqrt(2)/2) = 45 degrees

OpenStudy (anonymous):

ok so I do the same for my problem but I think Im getting the wrong answer. arccos((32√(41))/(205)) = 0.0312

jimthompson5910 (jim_thompson5910):

what did you get for the dot product

OpenStudy (anonymous):

32

OpenStudy (anonymous):

@jim_thompson5190?

jimthompson5910 (jim_thompson5910):

the length of vector u is sqrt(41) the length of vector v is 5 so this means cos(theta) = 32/(5*sqrt(41)) theta = arccos(32/(5*sqrt(41))) theta = 1.78991060824608 degrees

OpenStudy (anonymous):

That is radians, not degrees.

jimthompson5910 (jim_thompson5910):

no 1.78991060824608 is in degrees

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

That is weird.

jimthompson5910 (jim_thompson5910):

the two vectors are actually that close together (their lengths are)